OpenStudy (anonymous):

If ƒ(x ) = x 2 + 1 and g(x ) = 3x + 1, find ƒ(2) + g(3). 5/3 12 15

OpenStudy (anonymous):

I see my life save again!

OpenStudy (anonymous):

just find f(2) and then g3)

Nnesha (nnesha):

hahhaha :) :) :)

OpenStudy (anonymous):

f(2)=(2)*2+1 + 3(3)+1

OpenStudy (anonymous):

=4+1+9+1=15

Nnesha (nnesha):

@seancreed ok so this time other people gonna help you see you next time :) 0ok

OpenStudy (anonymous):

Thank you @Nnesha

Nnesha (nnesha):

no problem m'am :)